OpenStudy (anonymous):

Simplify the expression 4–3.

sam (.sam.):

4-3=1

OpenStudy (anonymous):

no its 4^-3

sam (.sam.):

\[4^{-3}=\frac{1}{4^3}=\frac{1}{64}\]

OpenStudy (anonymous):